# Few-Shot Learning on Graphs: from Meta-learning to Pre-training and Prompting

### v3

@@ Abstract
Graph representation learning, a critical step in graph-centric tasks, has seen significant advancements. Earlier techniques often operate in an end-to-end setting, where performance heavily relies on the availability of ample labeled data. This constraint has spurred the emergence of few-shot learning on graphs, where only a few task-specific labels are available for each task. Given the extensive literature in this field, this survey endeavors to synthesize recent developments, provide comparative insights, and identify future directions. We systematically categorize existing studies into three major families: meta-learning approaches, pre-training approaches, and hybrid approaches, with a finer-grained classification in each family to aid readers in their method selection process. Within each category, we analyze the relationships among these methods and compare their strengths and limitations. Finally, we outline prospective future directions for few-shot learning on graphs to catalyze continued innovation in this field.
1 Introduction
Data objects are often involved in complex interactions with each other, forming diverse network structures also known as graphs. The prevalence of graph-structured data gives rise to a surge in graph-centric tasks, such as node classification, link prediction, and graph classification. Recently, graph representation learning in a supervised setting Kipf and Welling (2017) has emerged as a prominent family of techniques for these graph-based tasks. The success of such techniques often depend on extensive labeled data for end-to-end training. However, acquiring labeled data is often challenging or expensive, limiting the performance of these supervised techniques Liu et al. (2023c).
Figure 1: Taxonomy of techniques for few-shot learning on graphs.
In light of this issue, few-shot learning on graphs Wang et al. (2020) has attracted significant interest in the field. Specifically, the number of labeled instances (termed as shots) in each task is notably small, representing a more practical setting in real-world applications. Few-shot learning approaches have been widely adopted in graph-based tasks, ranging from social network analysis to recommendation systems and molecular research. These approaches boil down to learning prior insights from data, often external to the target tasks, and subsequently adapting the prior knowledge to the downstream target tasks in a few-shot setup. Naturally, two key challenges arise. First, how can we learn a high-quality prior from graph data to provide useful references for downstream few-shot tasks? Second, how can we effectively and efficiently adapt the prior to downstream tasks?
To answer these challenges, this survey presents a concise yet systematic review of prevailing methods for few-shot learning on graphs. It aims to summarize recent developments, provide comparative insights, and identify future research avenues, contributing substantively to the few-shot learning and graph learning communities. For a systematic and comprehensive overview, we categorize the body of literature for few-shot learning on graphs into three major families: (1) Meta-learning approaches, (2) Pre-training approaches, and (3) Hybrid approaches, as shown in Fig. 1. Notably, both meta-learning Ding et al. (2020); Sheng et al. (2020) and pre-training Veličković et al. (2018); Qiu et al. (2020) represent two predominant research trends to learn priors from additional data, which are subsequently adapted for few-shot downstream tasks. Nevertheless, the two trends embrace distinct paradigms. On one hand, mainstream meta-learning approaches, such as Model Agnostic Meta-Learning (MAML) Finn et al. (2017) and Prototypical Networks (Protonets) Snell et al. (2017), derive a prior from a series of “meta-training” tasks (with task-specific labels) that mirror the downstream “meta-testing” tasks, assuming that all these tasks are drawn from an identical task distribution. They do not directly learn a model to serve as the prior. Rather, they aspire to learn a general adaptation process, also known as “learning-to-learn.” On the other hand, pre-training approaches utilize unlabeled data to optimize self-supervised pretext tasks, generally unrelated to the downstream tasks. They directly train a model as the prior, followed by an additional tuning phase to adapt the prior to specific downstream tasks. In the context of few-shot learning on graphs, meta-learning methods further exploit structure- or adaptation-based enhancements to cater to graph data. In contrast, pre-training approaches commonly train a graph encoder (i.e., the prior) in the pre-training phase, aiming to capture inherent characteristics of graph data, such as local connectivity or global topological patterns. Then, in the downstream phase, they either perform a full fine-tuning of the pre-trained graph encoder on each task, or employ a more parameter-efficient adaptation strategy. Finally, hybrid approaches integrating both paradigms have emerged recently, thereby synergistically harnessing the strengths of meta-learning and pre-training approaches. That is, they are able to exploit both unlabeled graph data and annotations on similar tasks, for a more seamless adaptation to downstream tasks.
Over the past few years, many efforts have been devoted to reviewing the few-shot learning literature from diverse perspectives Wang et al. (2020); Song et al. (2023); Wang et al. (2019a), but they do not cover few-shot learning on graphs. Recently, Zhang et al. Zhang et al. (2022) have discussed this field but limited its categorization to node-, edge- and graph-level problems in the meta-learning paradigm, lacking a comprehensive view of the technical approaches. Our survey endeavors to fill this gap by presenting a holistic overview of few-shot learning on graphs, proposing a fine-grained taxonomy of both classical and emerging methodologies. In summary, the contribution of our survey is threefold: 1) introduction of a novel taxonomy, enabling a systematic and concise overview of the literature in the field; 2) synthesis of recent progress within each branch of the taxonomy, coupled with comparative analyses of their respective strengths and weaknesses; 3) discussion of open challenges and promising directions for future research.
@@ 1 Introduction
The remainder of this survey is organized as follows. Sects. 2 and 3 delve into the meta-learning and pre-training paradigms, respectively, in the context of few-shot learning on graphs. Sect. 4 explores the hybrid approaches that integrate and benefit from both paradigms. Sect. 5 examines open challenges in the field and identifies potential avenues for future research. Finally, Sect. 6 concludes the survey.
2 Meta-learning Approaches

### v4

# A Survey of Few-Shot Learning on Graphs: from Meta-Learning to Pre-Training and Prompt Learning

@@ Abstract
Graph representation learning, a critical step in graph-centric tasks, has seen significant advancements. Earlier techniques often operate in an end-to-end setting, which heavily rely on the availability of ample labeled data. This constraint has spurred the emergence of few-shot learning on graphs, where only a few labels are available for each task. Given the extensive literature in this field, this survey endeavors to synthesize recent developments, provide comparative insights, and identify future directions. We systematically categorize existing studies based on two major taxonomies: (1) Problem taxonomy, which explores different types of data scarcity problems and their applications, and (2) Technique taxonomy, which details key strategies for addressing these data-scarce few-shot problems. The techniques can be broadly categorized into meta-learning, pre-training, and hybrid approaches, with a finer-grained classification in each category to aid readers in their method selection process. Within each category, we analyze the relationships among these methods and compare their strengths and limitations. Finally, we outline prospective directions for few-shot learning on graphs to catalyze continued innovation in this field. The website for this survey can be accessed by https://github.com/smufang/fewshotgraph.
Index Terms:
Graph mining, few-shot learning, meta-learning, pre-training, fine-tuning, prompting.
I Introduction
Data objects are often involved in complex interactions with each other, forming diverse network structures also known as graphs. The prevalence of graph-structured data has led to a growing demand in graph-centric tasks, such as node classification [1], link prediction [2], and graph classification [3]. Conventional graph analysis usually relies on feature engineering to exploit the structural information [4, 5, 6], necessitating significant manual effort. The emergence of graph embedding [7, 8, 9] opens up great opportunities for automated graph representation learning without requiring handcrafted features. Generally, these approaches encode various graph elements (e.g., nodes, edges or subgraphs) into low-dimensional vectors. By leveraging co-occurrence patterns among these elements, they aim to preserve graph structural information in a latent vector space [10]. Subsequently, graph neural networks (GNNs) [1, 11, 12] have integrated both structural information and node semantic features, achieving state-of-the-art performance in graph-centric tasks. Modern GNNs typically adopt a message-passing framework, where each node receives and aggregate features from neighboring nodes recursively. As a result, both structural and semantic information are incorporated into graph representations, empowering GNNs to attain state-of-the-art performance in various downstream tasks [13]. More recently, given the remarkable success of transformers in the fields of natural language processing (NLP) and computer vision (CV), transformer-based architectures [14, 15, 16, 17, 18] have also gained popularity in graph learning.
Figure 1: Overview of few-shot learning on graphs. (a–b) Different problem settings involving label or structure scarcity on graphs. (c–e) Different applications, such as cold-start recommendation on user-item graphs.
Figure 2: Overview of few-shot learning techniques on graphs.
I-A Few-Shot Learning on Graphs
Unfortunately, the effectiveness of prevailing graph representation learning techniques, including GNNs and graph transformers, depends on not only rich graph structures, but also extensive labeled data. However, data scarcity is common in real-world applications, significantly hindering their performance. Specifically, we observe data scarcity in two broad categories: (1) Label scarcity and (2) Structure scarcity. First, acquiring labeled data is often challenging or expensive [19], leading to label scarcity, especially for novel classes, as illustrated in Fig. 1(a). Second, graph structures may be sparse, especially for a large number of low-degree nodes with a limited number of neighbors, as shown in Fig. 1(b), resulting in structure scarcity.
In light of data scarcity, few-shot learning methods on graphs [20, 21] have attracted significant attention across various domains, such as social network analysis [22, 23], recommendation systems [24, 25], and molecular research [21, 26], as shown in Fig. 1(c,d,e). These approaches boil down to learning universal, task-independent prior knowledge from graphs, which is then adapted to tackle tasks with label scarcity or graphs with structure scarcity. Based on the specific strategies for learning prior knowledge and downstream adaptation, we further classify these techniques into three major families: (1) Meta-learning approaches, (2) Pre-training approaches, and (3) Hybrid approaches, as depicted in Fig. 2. Notably, both meta-learning [27, 28] and pre-training [29, 30] represent two predominant research trends for learning priors from additional data, which are then adapted for few-shot downstream tasks. However, these two trends follow distinct paradigms, as we further elaborate next.
Meta-learning on graphs. Mainstream meta-learning approaches, such as Model Agnostic Meta-Learning (MAML) [31] and Prototypical Networks (Protonets) [32], derive prior knowledge from a series of “meta-training” tasks (with task-specific labels) that mirror the downstream “meta-testing” tasks, assuming that all these tasks are independently drawn from an identical task distribution. Instead of directly learning a model, they aim to learn model initialization parameters or adaptation mechanisms that can be rapidly fine-tuned for new tasks—a concept known as “learning to learn.”
Meta-learning presents a compelling approach to few-shot learning on graphs. We introduce a novel taxonomy that focuses on the diverse ways in which graph data enhance existing meta-learning frameworks. Specifically, we identify two major forms of graph-enabled enhancement: structure-based and adaptation-based enhancements. On one hand, structure-based enhancement utilizes additional contexts provided by graphs [33, 34], categorized further into node-, edge-, and subgraph/graph-level contexts. These graph-grounded contexts provide complementary signals at various granularity levels, supplementing existing meta-learning methods for few-shot learning. On the other hand, adaptation-based enhancement focuses on leveraging graphs to enhance or refine the adaptation step [35, 36]. The adaptation step, crucial in many meta-learning frameworks, involves adapting the learned meta-knowledge to new tasks during testing. In particular, adaptation can be enhanced in various ways, including at the task, graph, and feature levels.
Pre-training on graphs. On the other hand, pre-training approaches utilize unlabeled data to optimize self-supervised pretext tasks, generally unrelated to the downstream tasks. They directly train a model as the prior, followed by an additional tuning phase to adapt the prior to specific downstream tasks [37, 38]. As pre-training emerges as a pivotal strategy for few-shot learning on graphs, we categorize these methods based on the specific techniques utilized during the pre-training and downstream adaptation phases.
In the pre-training phase, pre-training approaches commonly train a graph encoder (i.e., the prior), aiming to capture inherent characteristics of graph data, such as local connectivity [39, 30] or global topological patterns [40, 41]. These approaches adopt self-supervised pretext tasks that are either contrastive or generative. Specifically, contrastive strategies [29, 42] derive meaningful representations by contrasting positive and negative pairs of instances, whereas generative strategies [43, 39] train graph encoders and decoders to reconstruct graph elements, such as masked node features or dropped edges. Apart from pre-training methods that focus solely on graph data, recent developments [44, 45] have integrated language models into graph pre-training to exploit textual descriptions associated with graphs.
Following the pre-training phase, in the downstream adaptation phase, pre-training approaches either perform a fine-tuning of the pre-trained graph encoder on each task, or employ a more parameter-efficient adaptation strategy. Generally, fine-tuning [29, 30] updates the parameters in the pre-trained graph encoder and a task head based on labeled samples for each downstream task, thereby tailoring the pre-trained knowledge to each specific task. However, full fine-tuning can be expensive and often leads to overfitting in few-shot settings, where each task has only a few labeled samples. Alternatively, parameter-efficient adaptation freezes most or all of the parameters of the pre-trained model and tunes only a much smaller module. Specifically, parameter-efficient fine-tuning [46, 47] only updates a minimal subset of the pre-trained weights and the task head, or an additional lightweight module injected into the pre-trained model. In contrast, prompt-based learning [48, 19, 49] adopts a unified template that bridges pretext and downstream tasks, eliminating the need for any task head. Subsequently, it learns a task-specific prompt, consisting of only a small number of parameters, for each downstream task while freezing the pre-trained graph encoder.
Hybrid approaches. Finally, hybrid approaches that integrate both paradigms have emerged recently, thereby synergistically harnessing the strengths of meta-learning and pre-training approaches. These methods are able to exploit both unlabeled graph data and annotations from similar tasks, enabling more seamless adaptation to various downstream tasks.
I-B Differences with Existing Surveys and Our Contributions
Over the past few years, many efforts have been devoted to reviewing the few-shot learning literature from diverse perspectives [50, 51, 52], but they do not cover few-shot learning on graphs. Recently, Zhang et al. [53] have discussed this field but limited its categorization to node-, edge- and graph-level problems in the meta-learning paradigm, lacking a comprehensive review of the data scarcity problems and technical approaches. Besides, there are also some related reviews that have surveyed meta-learning on graphs [54], pre-training on graphs [55] and prompting on graphs [56, 57]. However, these surveys have predominantly focused on specific techniques for a single paradigm, tailored solely to the label scarcity problem, resulting in a relatively narrow coverage of the field.
Our survey endeavors to fill this gap by presenting a holistic overview of few-shot learning on graphs, proposing a multi-perspective and fine-grained taxonomy of both problems and methodologies. More specifically, we systematically categorize the literature into two novel taxonomies based on their problem settings and technical approaches, respectively. We divide the problem taxonomy into label scarcity and structure scarcity settings, while organizing the technique taxonomy into meta-learning, pre-training, and hybrid approaches. Each branch is further examined and subdivided into finer categories.
In summary, the contributions of our survey are threefold, as follows. (1) We introduce two taxonomies for few-shot learning on graphs in terms of problem settings and technical methodologies, respectively, providing a systematic and comprehensive overview of the literature in the field. (2) We synthesize recent progress within each branch of the taxonomies, offering comparative analyses of their respective strengths and weaknesses. (3) We discuss open challenges and outline promising directions for future research.
@@ I-B Differences with Existing Surveys and Our Contributions
The remainder of this survey is organized as follows. Section II introduces the relevant preliminaries and definitions used throughout the survey. Section III presents the taxonomy on data scarcity problems, discussing the issues of label scarcity and structure scarcity in the context of few-shot learning on graphs. Sections IVVI delve into the finer categories of the taxonomy on methodologies. More specifically, Sections IV and V cover the meta-learning and pre-training paradigms, respectively, while Section VI explores the hybrid approaches that combine both meta-learning and pre-training paradigms to leverage their respective advantages. Section VII outlines open challenges in the field and suggests potential directions for future research. Finally, Section VIII concludes the survey.
II Background
In this section, we introduce the preliminaries for graph data and graph representation learning. We summarize the frequently used notations in Table I.
TABLE I: Notations used in this paper.
Notation Description
$G$ A graph $G$.
$V$, $E$ Node set $\mathcal{V}$, edge set $E$.
$\mathbf{X}_{V}$, $\mathbf{X}_{E}$ The feature matrices of nodes and edges.
$T$, $R$ The set of node types and edge types.
$\phi$, $\varphi$ Type mapping functions for nodes/edges.
$\mathbf{x}_{v}$ The attribute of node $v$.
$\mathbf{h}_{v}$ ($\mathbf{h}^{l}_{v}$) The representation of node $v$ (in layer $l$).
$\mathcal{N}_{v}$ Neighbor set of node $v$.
$\theta$ The parameters set of a model.
$\omega$ The meta-knowledge.
$\mathcal{T}^{i}_{*}=(\mathcal{S}^{i}_{*},\mathcal{Q}^{i}_{*})$ Task $i$, support set $\mathcal{S}^{i}_{*}$, query set $\mathcal{Q}^{i}_{*}$.
$\mathcal{P}_{o}$ Positive samples of target instance $o$.
$\mathcal{N}_{o}$ Negative samples of target instance $o$.
$S_{v}$ Subgraph of ego-node $v$.
$\vec{s}_{v}$ Embedding of $S_{v}$.
II-A Graph Formalization
A graph is defined as $G=\{V,E,\mathbf{X}_{V},\mathbf{X}_{E},T,R,\phi,\varphi\}$, where $V$ is the set of nodes, $E$ is the set of edges, and $\mathbf{X}_{V}\in\mathbb{R}^{|V|\times d_{V}},\mathbf{X}_{E}\in\mathbb{R}^{|E|\times d_{E}}$ are the feature matrices of nodes and edges, respectively. Furthermore, $T$ and $R$ represent the set of node types and edge types, respectively. $\phi:V\rightarrow T$ is a mapping function that assigns each node $v\in V$ to its respective node type $\phi(v)\in T$, while $\varphi:E\rightarrow R$ is another mapping function that assigns each edge $e\in E$ to its corresponding edge type $\varphi(e)\in R$.
In different applications, graphs often manifest in diverse forms to preserve application-specific connectivity information. In its most basic form, a homogeneous graph $G_{\text{homo}}$ is a graph where $|T|=1$ and $|R|=1$, meaning that it contains only one type of node and edge. Conversely, a heterogeneous graph $G_{\text{hete}}$ [58], is a graph that satisfies $|T|>1$ or $|R|>1$. The rich semantic information conveyed by multiple node or edge types enables heterogeneous graphs to model many real-world scenarios [59, 60]. In particular, knowledge graphs[61] can be regarded as a form of heterogeneous graphs. Moreover, graphs often evolve over time in real-world applications. A dynamic graph ${G}_{\text{dyn}}=\{G^{(t)}:t\in\mathcal{T}\}$ models such scenarios, where $G^{(t)}$ represents the graph at time $t\in\mathcal{T}$ for some time domain $\mathcal{T}$, where the nodes, edges or their features can change over time. Additionally, the nodes or edges can be associated with multi-modal content, such as numerical, textual and visual features [44, 62], forming a multi-modal graph $G_{\text{mm}}$ where $\mathbf{X}_{V}$ or $\mathbf{X}_{E}$ is a feature tensor for the nodes or edges, respectively. In particular, if each node or edge is associated with textual content, we also call it a text-attributed graph.
II-B Graph Representation Learning
Graph representation learning [10] has emerged as a mainstream technique for graph analysis. It embeds graph structures into a low-dimensional space while preserving their inherent connectivity patterns. Formally, define $\mathbf{h}_{v}\in\mathbb{R}^{d}$ as the representation of node $v\in V$, formulated as:
$\mathbf{h}_{v}=f_{g}(v,G;\theta_{g}),$ (1)
where $f_{g}(\cdot;\theta_{g})$ parameterized by $\theta_{g}$ is the representation learning function, also known as the graph encoder, and $d$ is the dimension of the embedding vector $\mathbf{h}_{v}$.
Early graph representation learning approaches [7, 8, 9] usually exploit the co-occurrence relationships or proximity between nodes or substructures on graphs. More recently, graph neural networks (GNNs) capitalize on a message-passing framework, in which each node derives its representation by receiving and aggregating messages (i.e., features) from neighboring nodes, thus capturing both structural and semantic information. Multiple message-passing layers can be stacked such that, in the $l$-th layer, we obtain the representation $\mathbf{h}^{l}_{v}\in\mathbb{R}^{d_{l}}$ of node $v$ as
$\mathbf{h}^{l}_{v}=\textsc{Aggr}(\mathbf{h}^{l-1}_{v},\{\mathbf{h}^{l-1}_{u}:u\in\mathcal{N}_{v}\};\theta^{l}_{g}),$ (2)
where $\mathcal{N}_{v}$ is the set of neighbors of $v$, $\theta^{l}_{g}$ denotes the learnable parameters in the $l$-th layer, and $\textsc{Aggr}(\cdot)$ is a neighborhood aggregation function [63, 11]. Note that for the first layer, the input node embedding $\mathbf{h}^{0}_{v}$ is typically initialized with the node’s features from $\mathbf{X}_{V}$. For brevity, we denote the output node representations from the last layer as $\mathbf{h}_{v}$.
Finally, the whole-graph representation can be obtained through a readout operation:
$\mathbf{h}_{G}=\textsc{ReadOut}(\mathbf{h}_{v}:v\in V),$ (3)
which aggregates the node representations across the entire graph, such as sum pooling or mean pooling.
III Few-Shot Learning Problems on Graphs
Figure 3: Taxonomy of few-shot learning problems on graphs.
Few-shot learning on graphs has gained significant research interest due to the frequent data scarcity issues in real-world graphs. Based on the type of data scarcity, we categorize few-shot learning problems on graphs into two groups: label scarcity and structure scarcity. On one hand, similar to the challenges faced in NLP and CV [50, 51, 52], label scarcity, or the lack of labeled data, remains an important challenge in few-shot learning on graphs. On the other hand, unlike text and image data, graphs possess a non-euclidean topological structure. As a result, structure scarcity in graphs emerges as a second challenge that could adversely impact the learning of effective representations. Researchers have studied various problems related to the two challenges, and in this section, we aim to categorize the literature based on the taxonomy illustrated in Fig. 3.
III-A Label Scarcity Problem on Graphs
As acquiring labels is usually difficult or costly, label scarcity is a common problem in real-world applications. However, the performance of supervised methods heavily relies on a large amount of labeled data as supervision. Consequently, traditional supervised methods perform poorly when provided with limited labeled data, motivating few-shot learning approaches to address the label scarcity problem on graphs. We further categorize the label scarcity problem into class-based and instance-based label scarcity based on their respective class settings and target instances.
III-A1 Class-based Label Scarcity
Let $C$ denote the entire set of classes on a graph, which consists of two subsets: the base class set $C_{\text{base}}$ for model training, and the new class (also called novel or unseen class) set $C_{\text{new}}$ for testing, such that $C=C_{\text{base}}\cup C_{\text{new}}$ and $C_{\text{base}}\cap C_{\text{new}}=\emptyset$. Note that label scarcity could happen in either subsets or both, as follows.
Label scarcity in new classes. In this setting, the goal is to learn a prior using sufficient labeled data from the base classes $C_{\text{base}}$, and then transfer the prior knowledge to a novel task involving the new classes $C_{\text{new}}$. The labeled data for the new classes form the support set, while the unlabeled data from these new classes constitute the query set. When the support set contains exactly $K$ labeled samples for each of the $N$ classes from $C_{\text{new}}$, this setup is called an $N$-way $K$-shot problem. In particular, a few-shot problem refers to the scenario where $K$ is a small value. To address the few-shot problem for the new classes, researchers have developed meta-learning based approaches (see Sect. IV) to learn prior knowledge from labeled data in the base classes, and further transfer and adapt this knowledge to predict labels in the new classes [64, 65, 66, 36, 67, 68, 69, 70, 44, 49, 28]. However, the effectiveness of these approaches still heavily depends on the availability of abundant labeled data in a large number of base classes, which might also be challenging to obtain in real-world applications.
Label scarcity in base classes. This setting occurs when there is limited labeled data in $C_{\text{base}}$, resulting in insufficient prior knowledge to transfer and adapt to the new classes. More commonly, when there is no labeled data for the base classes or no base classes at all, meta-learning approaches become inapplicable. Therefore, researchers have turned to self-supervised methods to pre-train graph encoders without requiring any labeled data or base classes [43, 29, 42, 71, 39, 72, 30, 40, 73, 74, 75, 76, 77, 78, 79] (see Sect. V-A). The pre-trained model is subsequently adapted to novel downstream tasks involving new classes through fine-tuning the pre-trained weights (see Sect. V-B). The fine-tuning process typically requires a reasonable amount of labeled data and is usually not considered as a few-shot problem.
Label scarcity in both classes. This is a combination of the two settings above, where labeled data are limited in both $C_{\text{base}}$ and $C_{\text{new}}$. Similar to the label scarcity in base classes, self-supervised methods are often leveraged to first pre-train a graph encoder. However, with limited labeled data in the new classes as well, vanilla fine-tuning can be challenging in the few-shot context. Specifically, vanilla fine-tuning updates the full parameter set of the pre-trained model, which is not only expensive for large models but also prone to overfitting to the few-shot support set. Therefore, parameter-efficient adaptation methods [48, 19, 80, 46, 47, 81] are more suitable for these few-shot learning problems on graphs (see Sect. V-C).
III-A2 Instance-based Label Scarcity
An orthogonal categorization of label scarcity problems on graphs can be defined by the affected instances at node, edge or graph levels. We summarize the literature and their applications w.r.t. the instance-based categorization in Table II.
TABLE II: Summary of instance-based label scarcity learning on graphs and their applications.
Instance Application domain Literature
Node Academic network [64, 65, 66, 36, 67, 68, 69, 70, 44] [82, 83, 84, 48, 85, 86, 87, 49, 88]
Social network [64, 65, 36, 89, 69, 70, 48, 90, 49]
E-commerce network [65, 89, 68, 70, 49, 48]
Protein-protein interaction [91]
Traffic flow [92]
Edge Drug-drug interaction [93]
Protein multimer structure [94]
E-comm./academic network [95]
Knowledge graphs [96, 97, 98, 99, 100, 101, 102, 103, 104]
Graph Molecular graph [105, 106, 107, 108, 109, 110, 111]
Protein graph [112, 113, 94]
Social network [114, 49]
Node-level label scarcity. Nodes are fundamental components of graphs, making node-level tasks, such as node classification, highly prevalent. The lack of node labels leads to node-level label scarcity, where only a limited number of labeled nodes are available per class on the graph. In response to this challenge, a range of methods have been developed to address various node-centric applications. For example, in academic networks, where many authors or papers have unknown fields, researchers propose various strategies to classify authors or papers into their primary field, topic or venues with limited labels [64, 65, 66, 36, 67, 68, 69, 70, 44, 82, 83, 84, 48, 86, 87, 49, 88, 115, 116, 85]. In the realm of social networks, where interests of users or categories of organizations may be difficult or costly to obtain, previous studies [64, 65, 36, 89, 69, 70, 48, 90, 49, 115] propose effective solutions to overcome these obstacles. Moreover, with the growing popularity of e-commerce platforms, some recent research [65, 89, 68, 70, 49, 48, 115] tackles product categorization in e-commerce networks. Additionally, recent studies also explore node-level label scarcity in protein property prediction [91] and traffic speed prediction [92]. In a more challenging scenario, some studies address zero-shot node-level learning tasks in academic or e-commerce networks, where no labeled node is available at all [117, 118, 44, 82].
Edge-level label scarcity. Edges serve as explicit connectors between nodes in a graph. In knowledge graphs, various kinds of relation exist between entities. Hence, while nodes represent entities, edges are associated with labels that represent semantic relationships between entities [119, 120]. Likewise, in a drug-drug interaction network, edge labels indicate the nature of interaction between two drugs [121, 122]. However, in real-world applications, such labeled edges are often scarce: In knowledge graphs, it is impractical to exhaustively annotate all relationships, while in drug interactions, expensive laboratory setups may be needed. The edge-level scarcity motivates researchers to investigate solutions for learning with few-shot edge labels. In knowledge graphs, many works aim to predict the semantic relationship between entities [96, 97, 98, 99, 100, 101, 102, 103, 104], when novel relations are introduced with only one or a few labeled example available. In biochemical applications, few-shot solutions for drug interaction prediction [93] and protein multimer structure prediction [94] have been proposed. In e-commerce and academic networks, product-sharing prediction and co-authorship prediction have been explored in few-shot settings [95].
Graph-level label scarcity. Beyond predictions on individual nodes or edges, numerous important tasks are performed at the graph level, such as predicting properties or categories for subgraphs or whole graphs. Similar to the challenge at the node or edge level, we often face a shortage of labeled subgraph or graph samples. For example, several studies [114, 49] propose solutions for categorizing user communities in social networks. Moreover, few-shot approaches have been devised for property prediction for molecules [105, 106, 107, 108, 109, 110, 111] and proteins [112, 113, 94], where obtaining labels, such as solubility and permeability, often requires significant resources.
III-B Structure Scarcity Learning on Graphs
In contrast to label scarcity, which stems from the high cost or difficulty of obtaining annotations, structure scarcity arises from the inherent sparsity in graph topology, where many nodes are incident to very few edges. Since graph representation learning fundamentally relies on the richness of topological structures between nodes, structure scarcity presents a significant hurdle for effective learning. In the real world, structure scarcity in graphs is a common issue, which has attracted increasing research interest. We categorize the literature on structure scarcity into two subtypes: long-tailed distribution and cold-start learning problems, based on their goal and application scenarios as summarized in Table III. Specifically, the long-tailed problem addresses the challenge of learning from an imbalanced distribution, where a large number of nodes have few connections, while the cold-start problem focuses on learning representations for new nodes with no or very few connections.
TABLE III: Summary of structure scarcity learning on graphs and their applications.
Goal Application domain Literature
Long-tailed distribution Academic network [123, 124, 125, 126, 127, 128, 129]
Social network [123, 130, 124, 131, 132, 133, 126, 127, 128]
E-commerce network [133, 126, 134, 128, 129]
Protein-protein interaction [123]
Air traffic control [130]
Cold start Social network [135, 136, 137, 25, 138, 139]
E-commerce network [135, 25, 140]
Long-tailed distribution. Many datasets exhibit long-tailed distributions [141], where a large proportion of occurrences are concentrated in the low-frequency region, or the “tail”. In the context of graphs, only a small number of nodes, referred to as head nodes, are incident to a large number of edges (i.e., high-degree), while the majority of nodes, termed tail nodes, have very few edges (i.e., low-degree) [132]. Thus, tail nodes are surrounded by a small neighborhood with a scarcity of neighbor information, limiting effective learning of their representations. Formally, given a graph $G=(V,E)$, the set of head nodes is denoted by $V_{\text{head}}$, and the set of tail nodes by $V_{\text{tail}}$, such that $V_{\text{head}}\cup V_{\text{tail}}=V$ and $V_{\text{head}}\cap V_{\text{tail}}=\emptyset$. The separation of the two subsets is often determined by a predefined threshold on the node degree [124, 132].
Long-tailed distributions are common in real-world networks. For example, in academic networks, a small number of authors are highly prolific and linked to many papers, while the majority have only a few connections to papers [123, 124, 125, 126, 128, 129]. In social networks, a few celebrities have a vast number of followers, while most users have far fewer [123, 130, 124, 131, 132, 133, 126, 127, 128]. In e-commerce and recommendation systems [133, 126, 134, 128, 129], the disparity in the number of connections across users or items is also evident. The challenge of long-tailed distributions has also been studied in other graph-centric applications, ranging from protein-protein interaction [123] to air traffic control [130].
Cold-start learning. The cold-start problem in graphs arises when making predictions for new nodes, which often have no or very few existing edges. Conventionally, models are trained on existing graph structures where nodes have sufficient connections to other nodes, leading to poor performance for new nodes with few connections. This problem is prevalent in social networks and e-commerce systems, where new nodes are continually added. Specifically, in social networks, new users often begin with few interactions with existing users or content, motivating various solutions to the cold-start problem [135, 136, 137, 25, 138, 139]. Similarly, on commerce platforms, new users or items often lack interactions with other users or items during their initial phase [135, 25, 140].
Figure 4: Taxonomy of few-shot learning techniques on graphs.
IV Meta-Learning Techniques on Graphs
Meta-learning is an important family of few-shot learning techniques, represented in the first branch of our taxonomy in Fig. 4. Prevailing episodic meta-learning methods [142, 32, 31] are designed to learn a prior from base classes, which is transferable to new classes in downstream tasks. These methods typically assume an abundance of labeled data across a large number of base classes, while the downstream new classes have few labeled samples, i.e., the setting of label scarcity in new classes as discussed in Sect. III-A.
In this section, we first review standard meta-learning techniques. We then categorize and discuss literature that enhances the standard methods for few-shot learning on graphs.
IV-A Standard Meta-learning Techniques
In standard episodic meta-learning methods, such as Matching Networks (MN) [142], Model-Agnostic Meta-Learning (MAML) [31] and Prototpyical Networks (Protonets) [32], we first construct a series of meta-training tasks from the base classes to learn prior knowledge, which is then adapted to downstream few-shot meta-testing tasks involving the new classes. The concept is illustrated in Fig. 5, using few-shot node classification on a graph as an example. In this scenario, meta-training tasks are constructed using base classes with ample labels, whereas a meta-testing task involves new classes with only few-shot labels. Each task comprises a support set and a query set, which function as training and testing data, respectively, within each task [31]. The support set is always labeled, whereas the query set is only labeled in meta-training but is unlabeled for inference and prediction in meta-testing. The prior is optimized on the meta-training tasks to enable rapid, lightweight adaptation using the support set that can minimize the loss on the query set. Hence, for the prior to be transferable to meta-testing tasks, such meta-learners assume that the meta-training and -testing tasks are sampled from an i.i.d. task distribution.